# Self-Interacting Dark Matter with Mass Segregation: A Unified Explanation of Dwarf Cores and Small-Scale Lenses

### v2

@@ V Strong lensing perturbers
Figure 5 shows the evolution of the benchmark halos in the plane of projected inner mass (within 1 kpc) versus average projected slope $\gamma_{\rm 2D}$, with arrows indicating their trajectories over time. All three evolve toward the region favored by the SDSSJ0946+1006 perturber, with M08 passing through the 95% confidence contour at $t\approx 7.4$ Gyr. This demonstrates that mass segregation in two-component SIDM can drive certain halos into dense, lensing-relevant configurations even at modest cross sections. By contrast, in SIDM03, the same halos remain in the core-forming phase and evolve away from the dense regime required to match the perturber. To connect with the cosmological context, we overlay subhalos of the cluster host in SIDM2c, obtained using the parametric model. These subhalos extend to lower slopes and higher inner masses, with several candidates residing within the region favored by the SDSSJ0946+1006 perturber, supporting the viability of forming compact, dark structures in SIDM models with mass segregation.
VI Small-Scale Lenses and Lensing Cross Sections
@@
To quantify our framework’s capability to reconcile the GGSL anomaly, we perform cosmological simulations for CDM, SIDM2v, and SIDMx. The CDM simulation assumes a one-component model, while the latter two are detailed in Table 1. The resolution of these simulations is insufficient to make accurate predictions for dwarf halos. Therefore, we adopt a hybrid approach. We use simulated halo density profiles only for candidates with masses above $10^{13}~\rm M_{\odot}$ in the CDM simulation. These massive halos are matched across all three simulations, and only matched halos are retained, ensuring that the large-scale densities remain consistent with the respective cosmological runs. For lower-mass halos, we apply the parametric SIDM2c model to the corresponding CDM halos. We adopt the integral approach from Ref. Yang et al. (2024) to obtain predictions that incorporate the effects of accretion histories. This approach provides results that differ from the simulated SIDM2v and SIDMx models. For clarity, we refer to these hybrid models as Model-2v and Model-x, corresponding to cases where the massive halos are taken from the SIDM2v and SIDMx simulations, respectively. In effect, our study focuses on the impact of SIDM2c halos in the mass range $10^{10}$-$10^{12}~\rm M_{\odot}$ on GGSL. The integrated results are presented for two alternative models: one representing a scenario that remains safe from cluster SIDM constraints (Model-2v), and the other exhibiting extreme mass segregation (Model-x). We also incorporate the effect of baryons on halo density profiles using a parametric approach, obtaining results for the Model-2v+baryons and Model-x+baryons scenarios. For consistency with the CDM comparisons, the baryon potentials themselves are not included directly in the lensing analysis.
@@
To facilitate comparison with the GGSL measurment in Ref. Meneghetti et al. (2020), we fix the redshift of our host system at $z=0.439$ and compute lensing properties by including all subhalos within a $5~{\rm Mpc}/h$ sphere. The lensing potential and deflection angles are primarily modeled analytically using the parametric SIDM halo profile developed in Hou et al. (2025); Yang et al. (2024), which allows for efficient calculation of lensing observables across a wide range of halo masses and internal structures. See the Supplemental Material for details.
@@
Figure 6 compares the tangential (opaque) and radial (40% transparent) critical curves (top panels) and associated caustics (bottom panels) in the CDM, Model-2v, and Model-x simulations. We observe a substantial increase in the number and size of small-scale caustics in both two-component SIDM models compared to CDM.
@@
Quantitatively, we compute the fiducial GGSL cross section ($\sigma_{\rm GGSL}$, in unit of arcsec2) as the total area enclosed by secondary caustics (excluding the host caustic), within a $(-100^{\prime\prime},100^{\prime\prime})$ square region centered on the host halo. Following Meneghetti et al. Meneghetti et al. (2020), we apply a lower area threshold of $\pi(0.5^{\prime\prime})^{2}$. We also exclude any contribution with effective Einstein radius $\theta_{E}\equiv\sqrt{A_{c}/\pi}$ being larger than three arc-second, where $A_{c}$ is the area enclosed by a critical line. This cut prevents the statistical dominance of massive outliers, as they are physically distinct from dwarf halos and are not observed to produce $\theta_{E}>3^{\prime\prime}$ Meneghetti et al. (2022, 2023). We repeated the lensing analysis 11 times, corresponding to 11 different viewing angles. Each viewing angle was generated by consecutively rotating the system around the $x$- and $y$-axes in steps of $18^{\circ}$, i.e., $18^{\circ}$, $36^{\circ}$, $54^{\circ}$, $72^{\circ}$, and $90^{\circ}$ rotations for each axis, plus a case without rotation.
Table 2: GGSL cross section comparisons. The mean GGSL cross sections with their standard deviations and total number of contributing secondary caustics (aggregated over 11 analyses) are reported for different model scenarios. The ratio column computes the ratio of mean GGSL relative to the value in CDM, which illustrates the amount of excess. The GGSL cross section is computed considering secondary critical lines with $\theta_{E}<3^{\prime\prime}$.
@@
Table 2 summarizes the results of our GGSL analysis. The table presents, for each simulation, the mean GGSL cross section, its standard deviation, and the total number of contributing secondary caustics (aggregated over 11 analyses). The final column reports the ratio of the mean GGSL cross section relative to the CDM value, quantifying the amount of excess produced by two-component SIDM models. All values are computed considering only secondary critical lines with $\theta_{E}<3^{\prime\prime}$. The Model-2v and Model-x benchmarks correspond to host halos that are shallower and denser, respectively, than the CDM host halo. As expected, the GGSL cross section is higher in the denser Model-x case. In Model-2v, the GGSL cross section and the number of contributing caustics are both approximately double the CDM values. This enhancement increases to about 4.5 times the CDM value when baryons are included. The excess is systematically larger in Model-x, where the ratio rises to 7.35 without baryons and 13.7 with baryons. The level of GGSL excess produced by our benchmark models demonstrates their compelling capability to explain the observed GGSL excess, which is a factor of 3-6 above theoretical expectations.
VII Conclusion
@@
Future work could improve upon several aspects of this work, including the simulation resolution, the scan of model parameters, and the incorporation of baryonic physics. While controlled simulations are used to illustrate the effect of baryons, more dedicated simulations are required to map the full landscape of baryonic effects in SIDM2c halos. Moreover, the present study is limited to a single cluster system, which complicates direct comparison with observations.
@@
Despite these limitations, our current results highlight SIDM with mass segregation as a unified and testable framework for resolving multiple small-scale challenges. Looking ahead, our framework may also shed light on the formation of early massive galaxies and supermassive black holes, including the so-called “little red dots”, as mass segregation remains effective at high redshifts. Our results suggest that small-scale structure may already be revealing not only the strength but also the multi-component nature of dark matter.
Acknowledgements.
@@
We thank Moritz Fischer, Guoliang Li, Nan Li, Ran Li, Shihong Liao, Massimo Meneghetti, Priyamvada Natarajan, Huiyuan Wang, Hai-Bo Yu, and Yi-Ming Zhong for helpful discussion. The authors were supported in part by the National Key Research and Development Program of China (No. 2022YFF0503304), Natural Science Foundation of China under No. 12588101, Chinese Academy of Sciences (CAS), and the Project for Young Scientists in Basic Research of CAS (No. YSBR-092).
Appendix A SUPPLEMENTAL MATERIAL
Figure 7: Normalized density profiles for three $2\times 10^{10}~\rm M_{\odot}$ halos with median (blue), $1\sigma$ (green), and $2\sigma$ (orange) concentrations in the SIDM2c model, shown at increasing values of the gravothermal phase $\tau$. For comparison, the magenta curves show corresponding one-component SIDM profiles, and the initial NFW profile is shown in black. The red curves show the parametric model predictions for SIDM2c. The close match with the simulated curves demonstrates the quality of our calibration. The agreement among the three SIDM2c profiles (solid curves) in each panel demonstrates the conditioned universality of the model. Notably, the inner densities in the two-component case are consistently higher than in the one-component case apart from the late gravothermal phases ($\tau>1$), indicating that SIDM with mass segregation is more effective at enhancing the strong lensing signal.
Figure 8: Left: Normalized inner halo density evolution in SIDM2c for halos of $2\times 10^{10}~M_{\odot}$ and $10^{11}~M_{\odot}$ with various concentrations. The evolution time is normalized by the core collapse time calculated using Eq. 2 with $\sigma/m=14~\rm cm^{2}/g$. Middle: As in the left, but with the evolution time of all the simulations normalized to the first case in blue. Right: Model predictions (dash-dotted lines) versus simulation (shaded bands) for a $10^{11}~M_{\odot}$ halo ($1\sigma$ concentration, with baryons) at $t=0,3,6,9$ Gyr. Solid and dotted lines correspond to the halo and baryon profiles in the baryonic simulation, respectively. Dashed lines represent the halo profile in the DM-only simulation.
Appendix B A. A parametric model for the two-component model SIDM2c
B.1 Verification of the conditioned universality
@@
In the main text, we illustrated the existence of a conditioned universality in SIDM models with mass segregation by showing the rescaled halo density profile at a normalized gravothermal phase of $\tau\equiv t/t_{c}=0.6$. Here, we further explore these universal features in Fig. 7 by showing density profiles at additional gravothermal phases: $\tau=0.1,0.2,0.4,0.8,1.0,$ and $1.08$. For simplicity, we only show profiles for the three halos of mass $2\times 10^{10}~\rm M_{\odot}$ and compare them with the one-component SIDM results (dashed magenta curves) at the same gravothermal phases. The one-component results are obtained using a parametric model calibrated to the conduction fluid model from Ref. Hou et al. (2025).
@@
From Fig. 7, it is evident that the inner halo densities in the SIDM2c model are consistently higher than those in the one-component SIDM case across almost all gravothermal phases. Interestingly, in the late gravothermal phases ($\tau>1$), this difference diminishes, and the SIDM2c profiles become comparable to those of the one-component model. This result implies that the enhancement induced by SIDM2c does not stem from deeply core-collapsed halos, but rather from a broad population of halos across a range of gravothermal phases.
B.2 A parametric model for DM-only halos in SIDM2c
@@
$t_{c}$ $=$ $\frac{150}{C}\frac{1}{(\sigma/m)\rho_{s}r_{s}}\frac{1}{\sqrt{4\pi G\rho_{s}}},$ (2)
where we adopt $C=0.75$, and $\rho_{s}$ and $r_{s}$ are the NFW scale parameters for the CDM halo under consideration.
Figure 8 (left) illustrates the evolution of the inner halo density, where time is normalized by the core collapse time $t_{c}$ defined in Eq. 2. For each simulation, the inner density is computed at $r=0.05r_{s}$ using the NFW scale parameters $\rho_{s}$ and $r_{s}$, and normalized by $\rho_{s}$. Without the time rescale by $t_{c}$, the core collapse times for these halos range from 15 to 110 Gyr. A clear, unified evolutionary trend emerges after normalization, despite a roughly 30% uncertainty in $t_{c}$, which is comparable to the uncertainty in the one-component case.
Note that the conditioned universality in theory is independent of the $t_{c}$ formula. In Fig. 8 (middle), we rescale the evolution time in all the cases to the core collapse time of the first case shown. The agreement among the curves significantly improves over the left panel, hence future work can involve calibrating a more accurate expression for $t_{c}$ to reduce uncertainty in the parametric model.
@@
$\rho_{\rm PSIDM2c}(r|\rho_{s,0},r_{s,0},\tau)$ (3)
$=\frac{\rho_{s}}{\left[\left(\frac{r}{r_{s}}\right)^{4}+\left(\frac{r_{c}}{r_{s}}\right)^{4}\right]^{\frac{\gamma}{4}}\cdot\left[1+\left(\frac{r}{r_{s}}\right)^{\beta}\right]^{\frac{3-\gamma}{\beta}}},$
where $\rho_{s}=\rho_{s}(\rho_{s,0},\tau)$, $r_{s}=r_{s}(r_{s,0},\tau)$, $r_{c}=r_{c}(r_{s,0},\tau)$, $\beta=\beta(\tau)$ and $\gamma=\gamma(\tau)$ are functions of the phase parameter $\tau\in(0,1.08)$. The profile parameters as functions of $\tau$ and are fitted to be
$\rho_{s}(\tau)/\rho_{s,0}$ $=$ $\exp\left(-15.74\ln(3.748\tau+1)-0.4013+31.26\tau-15.58\tau^{2}+2.999\tau^{3}\right),$ (4)
$r_{s}(\tau)/r_{s,0}$ $=$ $1.116+7.066\tau+38.89\tau^{2}-217.3\tau^{3}+484.3\tau^{4}-557.2\tau^{5}+328.1\tau^{6}-77.84\tau^{7},$ (5)
$r_{c}(\tau)/r_{s,0}$ $=$ $0.5878\sqrt{\tau}-0.6073\tau+1.102\tau^{3/2}-1.384\tau^{2}-1.492\tau^{3}+5.162\tau^{4}-4.951\tau^{5}+1.634\tau^{6},$ (6)
$\beta(\tau)$ $=$ $1.128+9.554\tau-30.26\tau^{2}+124.9\tau^{3}-109.4\tau^{4}-181.9\tau^{5}+276.5\tau^{6},$ (7)
$\gamma(\tau)$ $=$ $0.2651+3.941\sqrt{\tau}+0.2936\tau-9.062\tau^{2}+14.35\tau^{3}-10.22\tau^{4}+2.785\tau^{5}$ (9)
@@
$V_{\rm max}(t)$ $=$ $V_{\rm max,CDM}(t_{f})+\int_{t_{f}}^{t}dt^{\prime}\frac{dV_{\rm max,CDM}(t^{\prime})}{dt^{\prime}}+\int_{t_{f}}^{t}\frac{dt^{\prime}}{t_{c}(t^{\prime})}\frac{dV_{\rm max,SIDM2c}(\tau^{\prime})}{d\tau^{\prime}}$
$R_{\rm max}(t)$ $=$ $R_{\rm max,CDM}(t_{f})+\int_{t_{f}}^{t}dt^{\prime}\frac{dR_{\rm max,CDM}(t^{\prime})}{dt^{\prime}}+\int_{t_{f}}^{t}\frac{dt^{\prime}}{t_{c}(t^{\prime})}\frac{dR_{\rm max,SIDM2c}(\tau^{\prime})}{d\tau^{\prime}},$ (14)
where $t_{f}$ is the halo formation time, chosen right after the halo’s exponential accretion phase. The derivatives $dV_{\rm max,SIDM2c}(\tau)/d\tau$ and $dR_{\rm max,SIDM2c}(\tau)/d\tau$ are obtained by differentiating the corresponding functions in Eq. 10, with the initial values $V_{\rm max,0}$ and $R_{\rm max,0}$ replaced by $V_{\rm max,CDM}(t^{\prime})$ and $R_{\rm max,CDM}(t^{\prime})$, respectively. After obtaining $V_{\rm max}(t)$ and $R_{\rm max}(t)$ for the corresponding $\tau$, the full SIDM2c density profile can be constructed by solving for the NFW parameters in Eq. 4. Further technical details are available in Ref. Yang et al. (2025b), which validates this approach using a population of halos from cosmological simulations. This method has also been applied to the SIDM Concerto simulation suite for tests across multiple mass scales Nadler et al. (2025).
B.3 Incorporating the influence of baryons

### v3

@@ Abstract
Fig. 5 shows the evolution of the benchmark halos in the plane of projected inner mass (within 1 kpc) versus average projected slope $\gamma_{\rm 2D}$, with arrows indicating their trajectories over time. All three evolve toward the region favored by the SDSSJ0946+1006 perturber, with M08 passing through the 95% confidence contour at $t\approx 7.4$ Gyr. This demonstrates that mass segregation in two-component SIDM can drive certain halos into dense, lensing-relevant configurations even at modest cross sections. By contrast, in SIDM03, the same halos remain in the core-forming phase and evolve away from the dense regime required to match the perturber. To connect with the cosmological context, we overlay subhalos of the cluster host in SIDM2c, obtained using the parametric model. These subhalos extend to lower slopes and higher inner masses, with several candidates residing within the region favored by the SDSSJ0946+1006 perturber, supporting the viability of forming compact, dark structures in SIDM models with mass segregation.
VI 6. Small-scale lenses and lensing cross sections
@@
To quantify our framework’s capability to reconcile the GGSL anomaly, we perform cosmological simulations for CDM, SIDM2v, and SIDMx. The CDM simulation assumes a one-component model, while the latter two are detailed in Table 1. The resolution of these simulations is insufficient to make accurate predictions for dwarf halos. Therefore, we adopt a hybrid approach. We use simulated halo density profiles only for candidates with masses above $10^{13}~\rm M_{\odot}$ in the CDM simulation. These massive halos are matched across all three simulations, and only matched halos are retained, ensuring that the large-scale densities remain consistent with the respective cosmological runs. For lower-mass halos, we apply the parametric SIDM2c model to the corresponding CDM halos. We adopt the integral approach from Ref. Yang et al. (2024) to obtain predictions that incorporate the effects of accretion histories. This approach provides results that differ from the simulated SIDM2v and SIDMx models. For clarity, we refer to these hybrid models as Model-2v and Model-x, corresponding to cases where the massive halos are taken from the SIDM2v and SIDMx simulations, respectively. In effect, our study focuses on the impact of SIDM2c halos in the mass range $10^{10}$$10^{12}~\rm M_{\odot}$ on GGSL. The integrated results are presented for two alternative models: one representing a scenario that remains safe from cluster SIDM constraints (Model-2v), and the other exhibiting extreme mass segregation (Model-x). We also incorporate the effect of baryons on halo density profiles using a parametric approach, obtaining results for the Model-2v+baryons and Model-x+baryons scenarios. For consistency with the CDM comparisons, the baryon potentials themselves are not included directly in the lensing analysis.
@@
To facilitate comparison with the GGSL measurement in Ref. Meneghetti et al. (2020), we fix the redshift of our host system at $z=0.439$ and compute lensing properties by including all subhalos within a $5~{\rm Mpc}/h$ sphere. The lensing potential and deflection angles are primarily modeled analytically using the parametric SIDM halo profile developed in Refs. Hou et al. (2025); Yang et al. (2024), which allows for efficient calculation of lensing observables across a wide range of halo masses and internal structures. See the Supplemental Material for details Balberg et al. (2002); Kormann et al. (1994); Moster et al. (2013); Carleton et al. (2019).
@@
Fig. 6 compares the tangential (opaque) and radial (40% transparent) critical curves (Fig. 6a, b, and c) and associated caustics (Fig. 6d, e, and f) in the CDM, Model-2v, and Model-x simulations. We observe a substantial increase in the number and size of small-scale caustics in both two-component SIDM models compared to CDM.
@@
Quantitatively, we compute the fiducial GGSL cross section ($\sigma_{\rm GGSL}$, in unit of arcsec2) as the total area enclosed by secondary caustics (excluding the host caustic), within a $(-100^{\prime\prime},100^{\prime\prime})$ square region centered on the host halo. Following Meneghetti et al. Meneghetti et al. (2020), we apply a lower area threshold of $\pi(0.5^{\prime\prime})^{2}$. We also exclude any contribution with effective Einstein radius $\theta_{E}\equiv\sqrt{A_{c}/\pi}$ being larger than three arcseconds, where $A_{c}$ is the area enclosed by a critical line. This cut prevents the statistical dominance of massive outliers, as they are physically distinct from dwarf halos and are not observed to produce $\theta_{E}>3^{\prime\prime}$ Meneghetti et al. (2022, 2023). We repeated the lensing analysis 11 times, corresponding to 11 different viewing angles. Each viewing angle was generated by consecutively rotating the system around the $x$- and $y$-axes in steps of $18^{\circ}$, i.e., $18^{\circ}$, $36^{\circ}$, $54^{\circ}$, $72^{\circ}$, and $90^{\circ}$ rotations for each axis, plus a case without rotation.
Table 2: GGSL cross section comparisons.
@@
The mean GGSL cross sections with their standard deviations and total number of contributing secondary caustics (aggregated over 11 analyses) are reported for different model scenarios. The ratio column computes the ratio of mean GGSL relative to the value in CDM, which illustrates the amount of excess. The GGSL cross section is computed considering secondary critical lines with $\theta_{E}<3^{\prime\prime}$.
Table 2 summarizes the results of our GGSL analysis. The table presents, for each simulation, the mean GGSL cross section, its standard deviation, and the total number of contributing secondary caustics (aggregated over 11 analyses). The final column reports the ratio of the mean GGSL cross section relative to the CDM value, quantifying the amount of excess produced by two-component SIDM models. All values are computed considering only secondary critical lines with $\theta_{E}<3^{\prime\prime}$. The Model-2v and Model-x benchmarks correspond to host halos that are shallower and denser, respectively, than the CDM host halo. As expected, the GGSL cross section is higher in the denser Model-x case. In Model-2v, the GGSL cross section and the number of contributing caustics are both approximately double the CDM values. This enhancement increases to about 4.5 times the CDM value when baryons are included. The excess is systematically larger in Model-x, where the ratio rises to 7.35 without baryons and 13.7 with baryons. The level of GGSL excess produced by our benchmark models demonstrates their potential to address the observed GGSL excess, which is a factor of 3-6 above theoretical expectations. Beyond the total GGSL cross section, we also compare the distribution of Einstein radii of the contributing secondary critical lines with observations. Additional details are provided in Appendices D and E.
VII 7. Conclusion
@@
Future work could improve upon several aspects of this work, including the simulation resolution, the scan of model parameters, and the incorporation of baryonic physics. While controlled simulations are used to illustrate the effect of baryons, more dedicated simulations are required to map the full landscape of baryonic effects in SIDM2c halos. In the present analysis, resolution limitations are mitigated using a parametric modeling approach, and the analysis is based on a single cluster zoom evaluated across 11 projections, which limits direct comparison with observations. Additional uncertainties can arise from host-dependent properties.
@@
Despite these limitations, the results support SIDM with mass segregation as a plausible framework for addressing multiple small-scale challenges. Looking ahead, our framework may also shed light on the formation of early massive galaxies and supermassive black holes, including the so-called “little red dots”, as mass segregation remains effective at high redshifts. Our results suggest that small-scale structure may already be revealing not only the strength but also the multi-component nature of dark matter.
Conflict of interest
The authors declare that they have no conflict of interest.
Acknowledgements
@@
We thank Moritz Fischer, Guoliang Li, Nan Li, Ran Li, Shihong Liao, Massimo Meneghetti, Priyamvada Natarajan, Huiyuan Wang, Hai-Bo Yu, and Yi-Ming Zhong for helpful discussion. The authors were supported in part by the National Key Research and Development Program of China (2022YFF0503304), Natural Science Foundation of China (12588101), Chinese Academy of Sciences (CAS), and the Project for Young Scientists in Basic Research of CAS (YSBR-092).
Author contributions
Daneng Yang and Yi-Zhong Fan proposed this study. Daneng Yang led the analysis and performed the simulations. Daneng Yang, Yi-Zhong Fan, and Yue-Lin Sming Tsai collaborated on writing the manuscript. Siyuan Hou contributed in part to the GGSL analysis. Yue-Lin Sming Tsai and Yi-Zhong Fan acquired funding for the analysis, particularly the simulations.
Appendix A A. A parametric model for the two-component model SIDM2c
A.1 Verification of the conditioned universality
@@
In the main text, we illustrated the existence of a conditioned universality in SIDM models with mass segregation by showing the rescaled halo density profile at a normalized gravothermal phase of $\tau\equiv t/t_{\rm c}=0.6$. Here, we further explore these universal features in Fig. S1 by showing density profiles at additional gravothermal phases: $\tau=0.1,0.2,0.4,0.8,1.0,$ and $1.08$. For simplicity, we only show profiles for the three halos of mass $2\times 10^{10}~\rm M_{\odot}$ and compare them with the one-component SIDM results (dashed magenta curves) at the same gravothermal phases. The one-component results are obtained using a parametric model calibrated to the conduction fluid model from Ref. Hou et al. (2025).
@@
From Fig. S1, it is evident that the inner halo densities in the SIDM2c model are consistently higher than those in the one-component SIDM case across almost all gravothermal phases. Interestingly, in the late gravothermal phases ($\tau>1$), this difference diminishes, and the SIDM2c profiles become comparable to those of the one-component model. This result implies that the enhancement induced by SIDM2c does not stem from deeply core-collapsed halos, but rather from a broad population of halos across a range of gravothermal phases.
A.2 A parametric model for DM-only halos in SIDM2c
@@
$t_{\rm c}$ $=$ $\frac{150}{C}\frac{1}{(\sigma/m)\rho_{\rm s}r_{\rm s}}\frac{1}{\sqrt{4\pi G\rho_{\rm s}}},$ (S1)
where we adopt $C=0.75$, and $\rho_{\rm s}$ and $r_{\rm s}$ are the NFW scale parameters for the CDM halo under consideration.
Fig. S2a illustrates the evolution of the inner halo density, where time is normalized by the core collapse time $t_{\rm c}$ defined in Eq. (S1). For each simulation, the inner density is computed at $r=0.05r_{\rm s}$ using the NFW scale parameters $\rho_{\rm s}$ and $r_{\rm s}$, and normalized by $\rho_{\rm s}$. Without the time rescale by $t_{\rm c}$, the core collapse times for these halos range from 15 to 110 Gyr. A clear, unified evolutionary trend emerges after normalization, despite a roughly 30% uncertainty in $t_{\rm c}$, which is comparable to the uncertainty in the one-component case.
Note that the conditioned universality in theory is independent of the $t_{\rm c}$ formula. In Fig. S2b, we rescale the evolution time in all the cases to the core collapse time of the first case shown. The agreement among the curves significantly improves over the left panel, hence future work can involve calibrating a more accurate expression for $t_{\rm c}$ to reduce uncertainty in the parametric model.
@@
$\rho_{\rm PSIDM2c}(r|\rho_{s,0},r_{s,0},\tau)$ (S2)
$=\frac{\rho_{\rm s}}{\left[\left(\frac{r}{r_{\rm s}}\right)^{4}+\left(\frac{r_{c}}{r_{\rm s}}\right)^{4}\right]^{\frac{\gamma}{4}}\cdot\left[1+\left(\frac{r}{r_{\rm s}}\right)^{\beta}\right]^{\frac{3-\gamma}{\beta}}},$
where $\rho_{\rm s}=\rho_{\rm s}(\rho_{s,0},\tau)$, $r_{\rm s}=r_{\rm s}(r_{s,0},\tau)$, $r_{c}=r_{c}(r_{s,0},\tau)$, $\beta=\beta(\tau)$ and $\gamma=\gamma(\tau)$ are functions of the phase parameter $\tau\in(0,1.08)$. The profile parameters as functions of $\tau$ and are fitted to be
$\rho_{\rm s}(\tau)/\rho_{\rm s,0}$ $=$ $\exp\left(-15.74\ln(3.748\tau+1)-0.4013+31.26\tau-15.58\tau^{2}+2.999\tau^{3}\right),$ (S3)
$r_{\rm s}(\tau)/r_{\rm s,0}$ $=$ $1.116+7.066\tau+38.89\tau^{2}-217.3\tau^{3}+484.3\tau^{4}-557.2\tau^{5}+328.1\tau^{6}-77.84\tau^{7},$ (S4)
$r_{c}(\tau)/r_{\rm s,0}$ $=$ $0.5878\sqrt{\tau}-0.6073\tau+1.102\tau^{3/2}-1.384\tau^{2}-1.492\tau^{3}+5.162\tau^{4}-4.951\tau^{5}+1.634\tau^{6},$ (S5)
$\beta(\tau)$ $=$ $1.128+9.554\tau-30.26\tau^{2}+124.9\tau^{3}-109.4\tau^{4}-181.9\tau^{5}+276.5\tau^{6},$ (S6)
$\gamma(\tau)$ $=$ $0.2651+3.941\sqrt{\tau}+0.2936\tau-9.062\tau^{2}+14.35\tau^{3}-10.22\tau^{4}+2.785\tau^{5}$ (S8)
@@
$V_{\rm max}(t)$ $=$ $V_{\rm max,CDM}(t_{f})+\int_{t_{f}}^{t}dt^{\prime}\frac{dV_{\rm max,CDM}(t^{\prime})}{dt^{\prime}}+\int_{t_{f}}^{t}\frac{dt^{\prime}}{t_{\rm c}(t^{\prime})}\frac{dV_{\rm max,SIDM2c}(\tau^{\prime})}{d\tau^{\prime}}$
$R_{\rm max}(t)$ $=$ $R_{\rm max,CDM}(t_{f})+\int_{t_{f}}^{t}dt^{\prime}\frac{dR_{\rm max,CDM}(t^{\prime})}{dt^{\prime}}+\int_{t_{f}}^{t}\frac{dt^{\prime}}{t_{\rm c}(t^{\prime})}\frac{dR_{\rm max,SIDM2c}(\tau^{\prime})}{d\tau^{\prime}},$ (S13)
where $t_{f}$ is the halo formation time, chosen right after the halo’s exponential accretion phase. The derivatives $dV_{\rm max,SIDM2c}(\tau)/d\tau$ and $dR_{\rm max,SIDM2c}(\tau)/d\tau$ are obtained by differentiating the corresponding functions in Eq. (S9), with the initial values $V_{\rm max,0}$ and $R_{\rm max,0}$ replaced by $V_{\rm max,CDM}(t^{\prime})$ and $R_{\rm max,CDM}(t^{\prime})$, respectively. After obtaining $V_{\rm max}(t)$ and $R_{\rm max}(t)$ for the corresponding $\tau$, the full SIDM2c density profile can be constructed by solving for the NFW parameters in Eq. (S3). Further technical details are available in Ref. Yang et al. (2025b), which validates this approach using a population of halos from cosmological simulations. This method has also been applied to the SIDM Concerto simulation suite for tests across multiple mass scales Nadler et al. (2025).
A.3 Incorporating the influence of baryons